A force of 27N is required to stretch a spring by 2.5cm. What is the spring constant
of the spring?
Answer #1

Solution) we have Force F = KX

K is spring constant

X is stretched distance

F = KX

K = F/X

Given F = 27 N

X = 2.5 cm = 2.5×10^(-2) m

K = F/X

K = (27)/(2.5×10^(-2))

K = 1080 N/m